OpenStudy (anonymous):

5/9v+w=z, for v show work please?

OpenStudy (compassionate):

\[\frac{ 5 }{ 9}v+w = z \rightarrow \frac{ 5 }{ 9}v+w-w = v - w \rightarrow \] \[\frac{ 5 }{ 9 }v = z - w \frac{ 9 }{ 5}\frac{ 5 }{ 9 }v = z - w \frac{ 9 }{ 5}\] v = z - w 9/5

OpenStudy (anonymous):

Thank You

OpenStudy (compassionate):

No problem